You notice that Coca-Cola has a stock price of $ 41.06 and EPS of $ 1.72. Its competitor PepsiCo has EPS of $ 3.98. But, Jones Soda, a small batch Seattle-based soda producer has a P/E ratio of 33.6 . Based on this information, what is one estimate of the value of a share of Pepsi Co stock? One share of PepsiCo stock is valued at $ (Round to the nearest cent.)
Answer #1

P/ E ratio = 41.06 / 1.72 = 23.872093

Value of Pepsi Co stock = Pepsi co EPS * P/E ratio = 3.98 * 23.872093 = 95.01